In the cylinder of a combustion engine, the total internal energy change from fuel combustion is -2537kJ for one "stroke". The cooling system surrounding the engine absorbs 947kJ as heat. Assuming the movement of the piston results in a volume expansion of 0.5 liters, what is the effective external pressure the piston expands against?
Answer #1

Change in internal energy = \Delta U = -2537 KJ

Heat is the released by the engine to the cooling system = q = -947KJ

Putting the values in the formula, \Delta U = -2537 KJ = q + P\times \Delta V = -947 + P \times 0.5

On solving, \Delta U = -2537 KJ = q + P\times \Delta V = -947 + P \times 0.5 \\ P = 3180 kPa